The beton package

Frank Jensen

1995/03/05

1 Introduction

The purpose of the beton package is to provide a setup so that documents using the package will automatically be typeset with the Concrete fonts designed by Don Knuth.

The beton package is based on Knuth’s macros for the book “Concrete Mathematics” [1]. Knuth’s macros can be found through anonymous ftp to labrea.

stanford.edu: look for the file gkpmac.tex in directory pub/tex/local/lib.

Also, the fonts can be found at that site: look in directory pub/tex/local/cm.

The Concrete fonts look best with a larger value for \baselineskip than the value normally used with the Computer Modern fonts. Therefore, the beton package makes some decisions about line spacing that override the values specified by the standard L^{A}TEX classes and packages. See Section ?? for further details.

In [1], the margin contains numerous humorous notes, so-called ‘graffiti’, contributed by the students that attended the course by Graham, Knuth, and Patashnik. These graffiti are typeset in a nine point slanted condensed version of the Concrete fonts. The current version of the beton package does not provide any commands to support such graffiti. However, such commands could easily be provided, e.g., by changing the \marginpar command.

The beton package deals with the text font issues of [1]; i.e., it does not make any decisions about page layout (such as the height and width of the text area), font size, numbering scheme (e.g., [1] uses the page number for the numbering of tables [but not figures, since the book doesn’t have any numbered illustrations]; use a separate package to get such a scheme), math fonts (use the euler package to get this distinctive part of the look of [1]), or the amount of space above and below displayed equations. However, the beton package does set up the following:

• Concrete fonts for text (Section ??).

• Increased value for \baselineskip (Section ??). This may not be appropriate for documents with short lines (e.g., two-column documents), so an option (standard-baselineskips) is provided to turn this feature off.

• An option (oldstyle-equation-numbers) to request ‘oldstyle’ numerals for equation numbers (Section ??).

• An option (concrete-math) to request ‘concrete’ math, if you insist (Section ??).^{1}

• Some commands to ‘fine tune’ the typography (Section ??).

Last, a note about the name of the present package: “beton” is the danish word for the ‘hard’ meaning of the english word “concrete” (and since there already exists a package of the name “concrete” by Frank Mittelbach and Rainer Schöpf, I chose “beton” for mnemonic reasons).

References

[1] R. L. Graham, D. E. Knuth, and O. Patashnik. Concrete Mathematics. Addison-Wesley, 1989.

1Footnote numbers are typeset in math mode, so if the math fonts are not properly set up, these numbers will come out wrong. The euler package and the concrete-math option will both provide a proper setup.